OpenStudy (simplixity):

On a road trip, a driver achieved an average speed of (48.0+A) km/h for the first 86.0 km and an average speed of (43.0-B) km/h for the remaining 54.0 km. What was her average speed (in km/h) for the entire trip? Round your final answer to three significant figures. A=16, B=8

OpenStudy (welshfella):

you need to use the formula average speed = distance / time taken Note you cannot average speeds because speeds are ratios.

OpenStudy (welshfella):

first calculate the time taken for both parts of the journey using t = d / s Then calculate the total os these 2 times then use s = d /t = (86+54) / t

OpenStudy (welshfella):

this will give you the average time for the entire trip

OpenStudy (welshfella):

time for first part of the trip = 86 / (48+16) h

OpenStudy (simplixity):

so I got about 1.34 for the first time? I think I'm doing this wrong..

OpenStudy (welshfella):

86 / 64 hors in the time for the first part

OpenStudy (simplixity):

Oh, so don't reduce. haha oops

OpenStudy (welshfella):

as a decimal this is 1.344 we need to correct it to 4 significant figures because the question want the answer correct to 3 sig figs.

OpenStudy (simplixity):

oh okay

OpenStudy (welshfella):

yes - you can reduce because the question wants answer as decimal

OpenStudy (welshfella):

second part, time = 54 / (43-8) = 54 / 35

OpenStudy (simplixity):

so the second time would be 53/35? which would be 1.514 ?

OpenStudy (welshfella):

= 1.543 i make it

OpenStudy (simplixity):

oh yeah 54, sorry!

OpenStudy (welshfella):

so total time = 1.344+ 1.543 = 2.887 total distance = 54 + 86 = 140 so average speed for whole journey is 140 / 2.887 km/h

OpenStudy (simplixity):

I got 48.49 km/h

OpenStudy (welshfella):

thats correct which to 3 sig figs is 48.5 km/h

OpenStudy (simplixity):

thank you so much!! :)

OpenStudy (welshfella):

yw